# Oka-1 manifolds: New examples and properties

### v2

@@ Abstract.
In this paper we investigate Oka-1 manifolds and Oka-1 maps, a class of complex manifolds and holomorphic maps recently introduced by Alarcón and Forstnerič. Oka-1 manifolds are characterised by the property that holomorphic maps from any open Riemann surface to the manifold satisfy the Runge approximation and Weierstrass interpolation conditions, while Oka-1 maps enjoy similar properties for liftings of maps from open Riemann surfaces in the absence of topological obstructions. We also formulate and study the algebraic version of the Oka-1 condition, called aOka-1. We show that it is a birational invariant for projective manifolds and holds for all uniformly rational projective manifolds. This gives a Runge approximation theorem for maps from compact Riemann surfaces to uniformly rational projective manifolds. Finally, we study a class of complex manifolds with an approximation property for holomorphic sprays of discs. This class lies between the smaller class of Oka manifolds and the bigger class of Oka-1 manifolds and has interesting functorial properties.
@@ 1. Introduction
Clearly, a complex manifold $X$ is Oka-1 if and only if every connected component of $X$ is such. By [2, Proposition 2.7], the conditions in Definition 1.1 are equivalent to the Runge approximation condition for any special pair of domains $K\subset K^{\prime}=K\cup D\subset R$, where $D$ is a closed disc attached to $K$ along a boundary arc, with jet interpolation at finitely many points of $K$. Every Oka manifold (see [16, 19]) is also an Oka-1 manifold; the converse is false in general, at least for noncompact manifolds.
@@
Alarcón and Forstnerič also introduced the following notion of an Oka-1 map, by analogy with the notion of an Oka map (for the latter, see [32, Sect. 16] and [12, Definition 7.4.7]).
@@
Algebraic Oka properties for morphisms from affine algebraic varieties of arbitrary dimension to algebraic manifolds were introduced and studied by Lárusson and Truong [34]. They showed in particular that no algebraic manifold which is compact or contains a rational curve satisfies the algebraic version of the basic Oka property, the approximation property, or the interpolation property [34, Theorem 2]. The most useful known algebraic Oka property is the relative Oka principle for morphisms from affine algebraic varieties to algebraically subelliptic manifolds; see [16, Definition 6.1] for this notion, due to Gromov [22], and [11, Theorem 3.1] and the more precise version in [16, Theorem 6.4] for the relevant result. It was recently shown by Kaliman and Zaidenberg [24, Theorem 0.1] that every algebraically subelliptic manifold is in fact algebraically elliptic, and this property is equivalent to several other algebraic Oka properties. (See also [16, Theorem 6.2 and Corollary 6.6].)
@@
Recall that a connected complex manifold $X$ is said to be rationally connected if any pair of points in $X$ can be joined by a rational curve $\mathbb{P}^{1}\to X$. Among many references for rationally connected projective manifolds, we refer to the papers by Kollár et al. [25, 27] and the monographs by Kollár [26] and Debarre [8]. It is shown in [27, Theorem 2.1] that several possible definitions of this class coincide. In particular, if every sufficiently general pair of points in $X$ can be connected by an irreducible rational curve, or by a chain of such curves, then $X$ is rationally connected. Also, if there is a morphism $f:\mathbb{P}^{1}\to X$ such that $f^{*}TX\to\mathbb{P}^{1}$ is an ample bundle, then $X$ is rationally connected. The class of projective rationally connected manifolds is birationally invariant. We have the following observation.
@@
In the present paper, we use different techniques to establish the aOka-1 property for two subclasses of the class of rationally connected projective manifolds: for rational manifolds (see Corollary 3.5), and for algebraically elliptic manifolds (see Theorem 4.1). Since the aOka-1 property is a birational invariant (see Corollary 3.4), these two results are summarised in Theorem 1.6. The theorem implies the following extension of Royden’s Runge approximation theorem [35, Theorem 10] for maps from compact Riemann surfaces to the Riemann sphere $\mathbb{P}^{1}$ to a much wider class of targets; in particular, to any projective space $\mathbb{P}^{n}$. The corollary is proved in Section 4.
@@
Namely, by a result of Stout [37, Theorem 8.1], every compact smoothly bounded domain in an open Riemann surface is biholomorphic to a domain in an affine Riemann surface, so an inductive application of the aOka-1 condition yields the ostensibly weaker but equivalent formulation of the Oka-1 property provided by [2, Proposition 2.7] (see the paragraph following Definition 1.1 above).
@@ 2. Oka-1 domains in Stein manifolds with the density property
The main result of this section (see Theorem 2.2) describes a large class of Oka-1 domains in Stein manifolds with the density property. It can be seen as an analogue of Kusakabe’s theorem [30, Theorem 1.2] that the complement of any compact holomorphically convex set in a Stein manifold with the density property is an Oka manifold.
@@
is an Oka-1 manifold. Furthermore, every holomorphic map $R\to X_{c}$  from an open Riemann surface $R$ can be approximated uniformly on a given compact set in $R$ and interpolated in finitely many points of $R$ by proper holomorphic maps $F:R\to X$ with $F(R)\subset X_{c}$.
@@
Pick a Riemannian distance function $\mathrm{dist}$ on $X$. Fix any $c_{0}\in\mathbb{R}_{+}$; we will show that the theorem holds for $X_{c_{0}}$. Let $A\subset K\subset R$, $\epsilon>0$, and $k\in\mathbb{N}$ be as in Definition 1.1, and let $f:R\to X$ be a continuous map which is holomorphic on an open neighbourhood $U\subset R$ of $K$ and satisfies $f(bK)\subset X_{c_{0}}$. Pick a smoothly bounded Runge domain $D$ in $R$ with $K\subset D\subset\bar{D}\subset U$ such that
@@
Since $\rho:X\to\mathbb{R}_{+}$ is a Morse exhaustion function whose Levi form has at least two positive eigenvalues at every point, [9, Theorem 1.1] and its proof (see in particular [9, Lemma 6.3]) furnishes a holomorphic map $\tilde{f}:\bar{D}\to X$ satisfying the following conditions.
@@
    $\sup_{p\in K}\mathrm{dist}(\tilde{f}(p),f(p))<\epsilon$.
@@
    $\rho(\tilde{f}(p))>\rho(f(p))-\epsilon$ for all $p\in\bar{D}$.
@@
    $\tilde{f}$ agrees with $f$ to order $k$ in every point of $A$.
  6. (f)
    $\tilde{f}$ is homotopic to $f|_{\bar{D}}$ by a homotopy of maps $f_{t}:\bar{D}\to X$ with $f_{0}=f|_{\bar{D}}$, $f_{1}=\tilde{f}$ satisfying $f_{t}(\overline{D\setminus K})\subset X_{0}$ for every $t\in[0,1]$.
(Note that condition (b) follows from (2.1) and (d) if $\epsilon>0$ is chosen small enough.) The assumptions imply that $\tilde{f}$ extends from $\bar{D}$ to a smooth map $\tilde{f}:R\to X$ satisfying the following two additional conditions:
@@
    The homotopy from $f|_{\bar{D}}$ to $\tilde{f}|_{\bar{D}}$ (see condition (e)) extends to a homotopy $f_{t}:R\to X$ $(t\in[0,1])$ satisfying $f_{t}(R\setminus D)\subset X_{c_{0}}$ for every $t\in[0,1]$.
Since the Stein manifold $X$ has the density property, the domain $O=X\setminus L$ in (2.2) is Oka by Kusakabe [30, Theorem 1.2]. Since $\tilde{f}(R\setminus D)\subset X_{c_{1}}\subset O$ (see (2.2) and (f)), it follows from [16, Theorem 1.3] that there is a holomorphic map $F:R\to X$ which approximates $\tilde{f}$ as closely as desired on $\bar{D}$, it agrees with $\tilde{f}$ (and hence with $f$) to order $k$ in the points of $A$, it satisfies $F(R\setminus D)\subset O$, and $F$ is homotopic to $f$ through maps $R\to X$ sending $R\setminus\mathring{K}$ to $X_{0}$. Thus, $F$ satisfies conditions (1) and (2) in Definition 1.1. In particular, if $f(K)\subset X_{c_{0}}$ then the map $F$, and the homotopy from $f$ to $F$, can be chosen to have range in $X_{c_{0}}$. This shows that $X_{c_{0}}$ is an Oka-1 manifold.
Furthermore, we can choose the map $F:R\to X_{c_{0}}$ as above to be a proper holomorphic immersion as a map to $X$. This follows from the main result of Andrist and Wold [5]. (See also [4, 13] for an extension of their result to maps from any Stein manifold $R$ satisfying $2\dim R\leq\dim X$.) However, by using the recently established fact that $X\setminus L$ is an Oka domain when $X$ is a Stein manifold with the density property and $L$ is a compact $\mathscr{O}(X)$-convex subset of $X$ (see Kusakabe [30, Theorem 1.2]), it is possible to give a simpler proof; cf. [17, Theorem 5.1]. The main idea is the following. We exhaust $R$ by an increasing sequence of compact $\mathscr{O}(R)$-convex subsets $K=K_{0}\subset K_{1}\subset\cdots\subset\bigcup_{i=0}^{\infty}K_{i}=R$. Likewise, we exhaust $X$ by compact $\mathscr{O}(X)$-convex subsets $L=L_{1}\subset L_{2}\subset\cdots\subset\bigcup_{i=1}^{\infty}L_{i}=X$, where $L$ is as above. (The sets $L_{i}$ are chosen as sublevel sets of a strongly plurisubharmonic Morse exhaustion function $\tau:X\to\mathbb{R}_{+}$.) We then inductively construct a sequence of continuous maps $f_{i}:R\to X_{c_{0}}$ $(i=0,1,\ldots)$ such that $f_{0}$ is the initial map, and for every $i=1,2,\ldots$ the map $f_{i}$ is holomorphic on $K_{i}$, it approximates $f_{i-1}$ as closely as desired on $K_{i-1}$ and agrees with it to order $k$ in the points of $A\subset K$, and it satisfies $f_{i}(R\setminus\mathring{K}_{i})\subset O_{i}=X\setminus L_{i}$. To obtain such a sequence, we alternately use [9, Theorem 1.1] and [16, Theorem 1.3], just like in the above construction of the map $\tilde{f}=f_{1}$ from the given initial map $f=f_{0}$. If the approximations are close enough then there is a limit holomorphic map $F=\lim_{i\to\infty}f_{i}:R\to X_{c_{0}}$, which is proper as a map to $X$ and satisfies the conditions in the theorem. ∎
@@
is such that the Levi form of its boundary $bL_{c}=\{\rho=c\}$ has at least one positive eigenvalue at every point (since the Levi form of $\rho$ has at least two positive eigenvalues at every point). Let us call a compact set $L\subset X$ with $\mathscr{C}^{2}$ boundary satisfying this property tangentially 1-convex.
@@ 3. Oka-1 is a bimeromorphic invariant
It was a long-standing open problem whether the Oka property is a bimeromorphic invariant, until Kusakabe proved that there are non-tame discrete subsets $A$ of $\mathbb{C}^{n}$, $n\geq 2$, for which $\mathbb{C}^{n}$ blown up at each point of $A$ is Brody volume-hyperbolic and hence not Oka [29, Example A.3]. Previously, it was shown that if $A$ is tame, then the blow-up is weakly subelliptic and hence Oka [12, Proposition 6.4.12]. It remains an open question whether an Oka manifold blown up at a single point is Oka.
On the algebraic side more is known. Kusakabe proved, by a reduction to [33, Theorem 1], that if an algebraic manifold $X$ satisfies the algebraic convex approximation property, aCAP, then so does the blow-up $B$ of $X$ along any algebraic submanifold (not necessarily connected); see [28, Corollary 4.3]. It follows that if $X$ is algebraically elliptic (see Section 4 for the definition and more information on this class), then $B$ is Oka. The optimal known geometric sufficient condition for a compact algebraic manifold to be algebraically elliptic is uniform rationality (see Arzhantsev et al. [6, Theorem 1.3]). If $X$ is uniformly rational, so is $B$ (see Gromov [22, 3.5E]). Hence, if $X$ is compact and uniformly rational, then $B$ is algebraically elliptic. If $X$ is of class $\mathcal{A}$ (see [12, Definition 6.4.4]), not necessarily compact, then $B$ is algebraically elliptic [33, Corollary 2]. Further, algebraic ellipticity implies strong dominability by affine spaces [33, Proposition 6], and if $X$ is dominable or strongly dominable, then so is $B$ [33, Theorem 9]. In view of all these results, it is reasonable to expect that an algebraically elliptic manifold blown up along an algebraic submanifold is algebraically elliptic, but this is yet to be proved.
@@
Next, assuming that $B$ is Oka-1, we take a continuous map $g:R\to X$ which is holomorphic on a neighbourhood of $K\cup A$. We wish to deform $g$ to a holomorphic map $R\to X$ which approximates $g$ on $K$ and agrees with $g$ to order $k(a)$ for each $a\in A$. Once we are able to lift $g$ to a continuous map $f:R\to B$, which is holomorphic on a neighbourhood of $K\cup A$, we simply apply the Oka-1 property of $B$ to obtain a suitable deformation of $f$ and then compose with $\pi$. Since the submanifold $Z$ of $X$ has codimension at least 2, by applying an arbitrarily small deformation to $g$ fixing every point $a\in A$ to order $k(a)$, we may assume that $g^{-1}(Z)\subset A$, and then the existence of a lifting $f$ is clear. ∎
@@
The corollary, combined with Royden’s Runge approximation theorem for maps from Riemann surfaces to $\mathbb{P}^{1}$ [35, Theorem 10], shows the following.
@@
Every compact rational manifold is Oka-1. In particular, the projective space $\mathbb{P}^{n}$ is aOka-1.
@@
By Royden’s theorem [35, Theorem 10], the Riemann sphere $\mathbb{P}^{1}$ is aOka-1. The required approximation and jet interpolation is precisely the content of his theorem. The required homotopy is then immediate, because an open Riemann surface has the homotopy type of a bouquet of circles and $\mathbb{P}^{1}$ is simply connected. Applying this result componentwise, it follows that every product $(\mathbb{P}^{1})^{n}$ of Riemann spheres is aOka-1. Since $(\mathbb{P}^{1})^{n}$ is a rational manifold, Corollary 3.4 concludes the proof. ∎

### v3

@@ Abstract.
In this paper we investigate Oka-1 manifolds and Oka-1 maps, a class of complex manifolds and holomorphic maps recently introduced by Alarcón and Forstnerič. Oka-1 manifolds are characterised by the property that holomorphic maps from any open Riemann surface to the manifold satisfy the Runge approximation and Weierstrass interpolation conditions, while Oka-1 maps enjoy similar properties for liftings of maps from open Riemann surfaces in the absence of topological obstructions. We also formulate and study the algebraic version of the Oka-1 condition, called aOka-1. We show that it is a birational invariant for compact algebraic manifolds and holds for all rational manifolds. This gives a Runge approximation theorem for maps from compact Riemann surfaces to uniformly rational projective manifolds. Finally, we study a class of complex manifolds with an approximation property for holomorphic sprays of discs. This class lies between the smaller class of Oka manifolds and the bigger class of Oka-1 manifolds and has interesting functorial properties.
@@ 1. Introduction
Clearly, a complex manifold $X$ is Oka-1 if and only if every connected component of $X$ is such. By [2, Proposition 2.7], the conditions in Definition 1.1 are equivalent to the Runge approximation condition for any special pair of domains $K\subset K^{\prime}=K\cup D\subset R$, where $D$ is a closed disc attached to $K$ along a boundary arc, with jet interpolation in finitely many points of $K$. Every Oka manifold (see [16, 19]) is also an Oka-1 manifold; the converse is false in general, at least for noncompact manifolds.
@@
Alarcón and Forstnerič also introduced the following notion of an Oka-1 map, by analogy with the notion of an Oka map (for the latter, see [33, Sect. 16] and [12, Definition 7.4.7]).
@@
Algebraic Oka properties for morphisms from affine algebraic varieties of arbitrary dimension to algebraic manifolds were introduced and studied by Lárusson and Truong [35]. They showed in particular that no algebraic manifold which is compact or contains a rational curve satisfies the algebraic version of the basic Oka property, the approximation property, or the interpolation property [35, Theorem 2]. The most useful known algebraic Oka property is the relative Oka principle for morphisms from affine algebraic varieties to algebraically subelliptic manifolds; see [16, Definition 6.1] for this notion, due to Gromov [22], and [11, Theorem 3.1] and the more precise version in [16, Theorem 6.4] for the relevant result. It was recently shown by Kaliman and Zaidenberg [24, Theorem 1.1] that every algebraically subelliptic manifold is in fact algebraically elliptic, and this property is equivalent to several other algebraic Oka properties. (See also [16, Theorem 6.2 and Corollary 6.6].)
@@
Recall that a connected complex manifold $X$ is said to be rationally connected if any pair of points in $X$ can be joined by a rational curve $\mathbb{P}^{1}\to X$. Among many references for rationally connected projective manifolds, we refer to the papers by Kollár et al. [26, 28] and the monographs by Kollár [27] and Debarre [8]. It is shown in [28, Theorem 2.1] that several possible definitions of this class coincide. In particular, if every sufficiently general pair of points in $X$ can be connected by an irreducible rational curve, or by a chain of such curves, then $X$ is rationally connected. Also, if there is a morphism $f:\mathbb{P}^{1}\to X$ such that $f^{*}TX\to\mathbb{P}^{1}$ is an ample bundle, then $X$ is rationally connected. The class of projective rationally connected manifolds is birationally invariant. We have the following observation.
@@
In the present paper, we use different techniques to establish the aOka-1 property for two subclasses of the class of rationally connected projective manifolds: for rational manifolds (see Corollary 3.5), and for algebraically elliptic manifolds (see Theorem 4.1). Since the aOka-1 property is a birational invariant (see Corollary 3.4), these two results are summarised in Theorem 1.6. The theorem implies the following extension of Royden’s Runge approximation theorem [36, Theorem 10] for maps from compact Riemann surfaces to the Riemann sphere $\mathbb{P}^{1}$ to a much wider class of targets; in particular, to any projective space $\mathbb{P}^{n}$. The corollary is proved in Section 4.
@@
Namely, by a result of Stout [38, Theorem 8.1], every compact smoothly bounded domain in an open Riemann surface is biholomorphic to a domain in an affine Riemann surface, so an inductive application of the aOka-1 condition yields the ostensibly weaker but equivalent formulation of the Oka-1 property in [2, Proposition 2.7] (see the paragraph following Definition 1.1 above).
Remark 1.10.
By [6, Corollary 1.8], a compact algebraic manifold $X$ of dimension $n$ is unirational if and only if there is a surjective morphism $\mathbb{C}^{n}\to X$. Then $X$ is densely dominable by $\mathbb{C}^{n}$, so $X$ is Oka-1 by [2, Theorem 2.2]. (We thank an anonymous referee for pointing this out to us.) Whether $X$ is in fact Oka or aOka-1 is unknown.
@@ 2. Oka-1 domains in Stein manifolds with the density property
The main result of this section (see Theorem 2.2) describes a large class of Oka-1 domains in Stein manifolds with the density property. It can be seen as an analogue of Kusakabe’s theorem [31, Theorem 1.2] that the complement of any compact holomorphically convex set in a Stein manifold with the density property is an Oka manifold.
@@
is an Oka-1 manifold. Furthermore, given a compact Runge set $K$ in an open Riemann surface $R$, a finite set $A\subset K$, and a continuous map $f:R\to X$ which is holomorphic on an open neighbourhood $U\subset R$ of $K$ and satisfies $f(R\setminus\mathring{K})\subset X_{c}$, there is a proper holomorphic map $F:R\to X$ which approximates $f$ as closely as desired uniformly on $K$, agrees with $f$ to a given finite order in the points of $A$, and satisfies $F(R\setminus\mathring{K})\subset X_{c}$.
@@
The first claim in the theorem, that $X_{c}$ is an Oka-1 domain, follows from the second part of the theorem and [2, Proposition 2.7], which gives a criterion for a complex manifold to be Oka-1 by a simpler approximation condition and with interpolation in finitely many points.
We now prove the second part of the theorem. We may assume that $R$ and $X$ are connected. It clearly suffices to prove the result for a domain $X_{c_{0}}=\{\rho>c_{0}\}$ where $c_{0}\in\mathbb{R}_{+}$ is a regular value of $\rho$. Let $A\subset K\subset U\subset R$ and $f:R\to X$ be as in the theorem with $c=c_{0}$. Pick a smoothly bounded Runge domain $D$ in $R$ with $K\subset D\subset\bar{D}\subset U$ such that
@@
Since $\rho:X\to\mathbb{R}_{+}$ is a Morse exhaustion function whose Levi form has at least two positive eigenvalues at every point, [9, Theorem 1.1] and its proof (see in particular [9, Lemma 6.3]) furnishes a holomorphic map $\tilde{f}:\bar{D}\to X$ satisfying the following conditions for any given pair of numbers $\epsilon>0$ and $k=1,2,\ldots$.
@@
    $\sup_{p\in K}\mathrm{dist}_{X}(\tilde{f}(p),f(p))<\epsilon$.
@@
    $\tilde{f}$ agrees with $f$ to order $k$ in every point of $A$.
@@
    $\tilde{f}$ is homotopic to $f|_{\bar{D}}$ by a homotopy of maps $f_{t}:\bar{D}\to X$ with $f_{0}=f|_{\bar{D}}$, $f_{1}=\tilde{f}$ satisfying $f_{t}(\overline{D\setminus K})\subset X_{c_{0}}$ for every $t\in[0,1]$.
The assumption on the Levi form of $\rho$ implies that it has Morse index at most $22\dim_{\mathbb{C}}X-$ 2 at every critical point. Hence, for any pair of regular values $c<c^{\prime}$ of $\rho$ the set $X_{c}=\{\rho>c\}$ is topologically obtained from $X_{c^{\prime}}$ by attaching handles of dimension at least $2$. It follows that the inclusion $X_{c^{\prime}}\hookrightarrow X_{c}$ induces a bijection of the path components and the relative fundamental group $\pi_{1}(X_{c},X_{c^{\prime}})$ vanishes. Since the pair $(R,D)$ is homotopy equivalent to a relative $1$-dimensional CW complex, we infer that $\tilde{f}$ extends from $\bar{D}$ to a continuous map $\tilde{f}:R\to X$ satisfying the following two additional conditions:
@@
    The homotopy $f_{t}:\bar{D}\to X$ in condition (e) extends to a homotopy $f_{t}:R\to X$ $(t\in[0,1])$ satisfying $f_{0}=f$, $f_{1}=\tilde{f}$, and $f_{t}(R\setminus D)\subset X_{c_{0}}$ for every $t\in[0,1]$.
Since the Stein manifold $X$ is assumed to have the density property, the domain $O=X\setminus L$ in (2.2) is Oka by Kusakabe [31, Theorem 1.2]. From the condition $\tilde{f}(R\setminus D)\subset X_{c_{1}}\subset O$ (see (2.2) and (f)) and [16, Theorem 1.3] it follows that there is a holomorphic map $F:R\to X$ which approximates $\tilde{f}$ as closely as desired on $\bar{D}$, it agrees with $\tilde{f}$ (and hence with $f$) to order $k$ in the points of $A$, it satisfies $F(R\setminus D)\subset O$, and $F$ is homotopic to $f$ through maps $R\to X$ sending $R\setminus\mathring{K}$ to $X_{0}$. Thus, $F$ satisfies conditions (1) and (2) in Definition 1.1. In particular, if $f(K)\subset X_{c_{0}}$ then the map $F$, and the homotopy from $f$ to $F$, can be chosen to have range in $X_{c_{0}}$. By [2, Proposition 2.7] this shows that $X_{c_{0}}$ is an Oka-1 manifold.
Furthermore, we can choose the map $F:R\to X$ as above to be a proper holomorphic immersion (embedding if $\dim X\geq 3$) provided the jet interpolation conditions allow it. This follows from the main result of Andrist and Wold [5]. (See also [4, 13] for an extension of their result to maps from any Stein manifold $R$ satisfying $2\dim R\leq\dim X$.) By using the recently established fact that $X\setminus L$ is an Oka domain when $X$ is a Stein manifold with the density property and $L$ is a compact $\mathscr{O}(X)$-convex subset of $X$ (see Kusakabe [31, Theorem 1.2]), a simpler proof is possible; cf. [17, Theorem 5.1]. The main idea is the following. We exhaust $R$ by an increasing sequence of compact $\mathscr{O}(R)$-convex subsets $K=K_{0}\subset K_{1}\subset\cdots\subset\bigcup_{i=0}^{\infty}K_{i}=R$. Likewise, we exhaust $X$ by compact $\mathscr{O}(X)$-convex subsets $L=L_{1}\subset L_{2}\subset\cdots\subset\bigcup_{i=1}^{\infty}L_{i}=X$, where $L$ is as above. (The sets $L_{i}$ are chosen as sublevel sets of a strongly plurisubharmonic Morse exhaustion function $\tau:X\to\mathbb{R}_{+}$.) We then inductively construct a sequence of continuous maps $f_{i}:R\to X$ $(i=0,1,\ldots)$ such that $f_{0}$ is the initial map, and for every $i=1,2,\ldots$ the map $f_{i}$ is holomorphic on $K_{i}$, it approximates $f_{i-1}$ as closely as desired on $K_{i-1}$ and agrees with it to order $k$ in the points of $A\subset K$, and it satisfies $f_{i}(R\setminus\mathring{K}_{i})\subset O_{i}=X\setminus L_{i}$. To obtain such a sequence, we alternately use [9, Theorem 1.1] and [16, Theorem 1.3], just like in the above construction of the map $\tilde{f}=f_{1}$ from the given initial map $f=f_{0}$. If the approximations are close enough then there is a limit holomorphic map $F=\lim_{i\to\infty}f_{i}:R\to X$, which is proper as a map to $X$ and satisfies the conditions in the theorem. ∎
@@
is such that the Levi form of its boundary $bL_{c}=\{\rho=c\}$ has at least one positive eigenvalue at every point (since the Levi form of $\rho$ has at least two positive eigenvalues at every point). Let us call a compact set $L\subset X$ with $\mathscr{C}^{2}$ boundary satisfying this property tangentially 1-convex. Note that every such set $L$ is strongly pseudoconvex if $\dim X=2$, but this fails if $\dim X\geq 3$. A specific example is $L=\{\rho\leq c\}\subset\mathbb{C}^{3}$ where $\rho(z_{1},z_{2},z_{3})=|z_{1}|^{2}+|z_{2}|^{2}-|z_{3}|^{2}+|z_{3}|^{4}$ and $c>0$ is a regular value of $\rho$. Indeed, $\rho$ is an exhaustion function on $\mathbb{C}^{3}$ whose Levi form has eigenvalues $1,1,-1+4|z_{3}|^{2}$ at each point, so $L$ is tangentially 1-convex. However, the signature of the Levi form of $bL$ along the circle $\{|z_{1}|^{2}=c,\ z_{2}=z_{3}=0\}\subset bL$ is $(1,-1)$, so $L$ is not strongly pseudoconvex.
@@ 3. Oka-1 is a bimeromorphic invariant
It was a long-standing open problem whether the Oka property is a bimeromorphic invariant, until Kusakabe proved that there are non-tame discrete subsets $A$ of $\mathbb{C}^{n}$, $n\geq 2$, for which $\mathbb{C}^{n}$ blown up at each point of $A$ is Brody volume-hyperbolic and hence not Oka [30, Example A.3]. Previously, it was shown that if $A$ is tame, then the blow-up is weakly subelliptic and hence Oka [12, Proposition 6.4.12]. It remains an open question whether an Oka manifold blown up at a single point is Oka.
On the algebraic side more is known. Kusakabe proved, by a reduction to [34, Theorem 1], that if an algebraic manifold $X$ satisfies the algebraic convex approximation property, aCAP, then so does the blow-up $B$ of $X$ along any algebraic submanifold (not necessarily connected); see [29, Corollary 4.3]. It follows that if $X$ is algebraically elliptic (see Section 4 for the definition and more information on this class), then $B$ is Oka. The optimal known geometric sufficient condition for a compact algebraic manifold to be algebraically elliptic is uniform rationality (see Arzhantsev et al. [6, Theorem 1.3]). A compact algebraic manifold $X$ is called uniformly rational if every point of $X$ admits a Zariski open neighbourhood isomorphic to a Zariski open set in $\mathbb{C}^{n}$ with $n=\dim X$. If $X$ is uniformly rational, so is the blow-up $B$ of $X$ along any algebraic submanifold (see Gromov [22, 3.5E]). Hence, if $X$ is compact and uniformly rational, then $B$ is algebraically elliptic. If $X$ is of class $\mathcal{A}$ (see [12, Definition 6.4.4]), not necessarily compact, then $B$ is algebraically elliptic [34, Corollary 2]. Further, algebraic ellipticity implies strong dominability by affine spaces [34, Proposition 6], and if $X$ is dominable or strongly dominable, then so is $B$ [34, Theorem 9]. In view of all these results, it is reasonable to expect that an algebraically elliptic manifold blown up along an algebraic submanifold is algebraically elliptic, but this is yet to be proved.
@@
Next, assuming that $B$ is Oka-1, we take a continuous map $g:R\to X$ which is holomorphic on a neighbourhood of $K\cup A$. We wish to deform $g$ to a holomorphic map $R\to X$ which approximates $g$ on $K$ and agrees with $g$ to order $k(a)$ for each $a\in A$. Once we are able to lift $g$ to a continuous map $f:R\to B$, which is holomorphic on a neighbourhood of $K\cup A$, we simply apply the Oka-1 property of $B$ to obtain a suitable deformation of $f$ and then compose with $\pi$. Since the submanifold $Z$ of $X$ has codimension at least $2>\dim R$, by applying an arbitrarily small deformation to $g$ fixing every point $a\in A$ to order $k(a)$, we may assume that $g^{-1}(Z)\subset A$, and then the existence of a lifting $f$ is clear. ∎
@@
The corollary, combined with Royden’s Runge approximation theorem for maps from Riemann surfaces to $\mathbb{P}^{1}$ [36, Theorem 10], shows the following.
@@
Every compact rational manifold is aOka-1. In particular, the projective space $\mathbb{P}^{n}$ is aOka-1.
@@
By Royden’s theorem [36, Theorem 10], the Riemann sphere $\mathbb{P}^{1}$ is aOka-1. The required approximation and jet interpolation is precisely the content of his theorem. The required homotopy is then immediate, because an open Riemann surface has the homotopy type of a bouquet of circles and $\mathbb{P}^{1}$ is simply connected. Applying this result componentwise, it follows that every product $(\mathbb{P}^{1})^{n}$ of Riemann spheres is aOka-1. Since $(\mathbb{P}^{1})^{n}$ is a rational manifold, Corollary 3.4 concludes the proof. ∎